# Displacement-Squeeze receiver for BPSK displaced squeezed vacuum states surpassing the coherent-states Helstrom bound under imperfect conditions

### v1

@@ Abstract
We propose a displacement-squeeze receiver (DSR) for discriminating BPSK displaced squeezed vacuum states (S-BPSK). The receiver applies a displacement followed by a squeezing operation with the squeezing axis rotated by $\frac{\pi}{2}$, and performs photon-number-resolving detection with a MAP threshold decision. This processing effectively increases the distinguishability of the input states by elongating their distance in phase space and reducing their population overlap in Fock basis. We show that for all signal energy N, $P_{\text{err}}^{\text{DSR}}\in\left[P_{\text{HB}}^{\text{DSS}},2P_{\text{HB}}^{\text{DSS}}\right]$, under equal priors and ideal condition. In the low-energy regime, DSR beats the S-BPSK SQL at $N\approx 0.3$ and drops below the coherent-state BPSK (C-BPSK) Helstrom bound at $N\approx 0.4$, reaching $P_{\text{err}}^{\text{DSR}}<1\%$ near $N\approx 0.6$. Finally, we quantify performance under non-unit efficiency and dark counts, phase diffusion, and receiver thermal noise, with MAP threshold adaptation providing robustness across these nonidealities.
@@ I Introduction
As an important topic in the field of quantum information, quantum-states discrimination directly affects the performance of quantum communication systems. To date, we have consistently regarded coherent states (CSs) as carriers for quantum communication and have proposed numerous excellent discrimination schemes, such as the Kennedy receiver[15], the optimized displacement receiver[24], multi-stage receiver[25, 14, 5, 18], the conditional pulse nulling receiver[32], and hybrid receiver [17, 1, 21, 33]. This is primarily because coherent states retain their coherence after passing through lossy bosonic channels, demonstrating that they can maintain state purity in most optical transmission media, including but not limited to free space and optical fibers.[8, 27, 16]. However, coherent states have inherent discrimination limitations: any two coherent states are non-orthogonal and exhibit overlap in phase space. Since the Helstrom Bound (HB) is closely related to the degree of overlap between quantum states[10, 11], coherent states have a theoretical lower bound on discrimination error probability.
In recent years, quantum-state discrimination schemes based on non-Poissonian photon statistics have attracted widespread attention[22, 13, 23, 6, 3, 26]. For example, Giovanni Chesi and Stefano Olivares proposed using squeezed vacuum states ($\hat{S}(r)\left|0\right\rangle$ where $\hat{S}(r),r\in\mathbf{R}_{+}$ is squeeze operator and $\left|0\right\rangle$ is vacuum state) as information carriers to construct a binary communication channel[6]. By applying a displacement operation $\hat{D}(\pm\alpha)$ to the squeezed vacuum state, two phase-opposite displaced squeezed vacuum states (DSSs) $\left|\pm\alpha,r\right\rangle$ can be generated. Then they encoded BPSK symbols into two phase-opposite DSSs (S-BPSK):
@@
This scheme significantly reduces the overlap between the carrier states by squeezing its position quadrature $\hat{X}$. Theoretical studies show that when the communication channel energy $N$ is fixed, not only the HB for S-BSPK can be significantly lower than that for BPSK coherent states (C-BPSK), but the error probability of homodyne detection (Standard Quantum Limit, SQL) for S-BSPK can even surpass the HB for C-BSPK at a low mean photon number of $N\approx 0.659$ [6].
Unlike the work of Giovanni Chesi and Stefano Olivares, this paper proposes a quantum-enhanced receiver for discriminating the S-BPSK signal, which is called as the Displacement-Squeeze Receiver (DSR). The DSR combines Gaussian operations (displacement and squeeze) and photon-number-resolving detection to approach the Helstrom bound for S-BPSK within a constant factor, while outperforming coherent-state receivers in the low-energy regime. Theoretical analysis shows that, for the same input mean photon number, the DSR can achieve an error probability lower than the Helstrom bound for C-BPSK channels in the low-photon-number regime. For example, when $N\approx 0.6$ the error probability falls below 1% in our simulations.
The structure of this paper is as follows: Section II reviews the fundamental theory of S-BPSK discrimination; Section III provides a detailed analysis of the principle and performance of the DSR; Section IV explores the performance of DSR under non-ideal conditions, including non-ideal PNR detectors (quantum efficiency $\eta$ and dark count rate $\nu$), phase diffusion noise in the channel, and receiver thermal noise; finally, the paper is summarized, and potential application scenarios are discussed.
@@ II Channel model and performance benchmarks
In this section we summarize the signal model and the fundamental performance limits for binary phase-shift keyed (BPSK) communication based on displaced squeezed vacuum states, and we fix the notation used in the rest of the paper.
For binary discrimination between quantum states $\left|\psi_{0}\right\rangle$ and $\left|\psi_{1}\right\rangle$ with prior probabilities $P_{0}=P_{1}=\frac{1}{2}$, the minimum error probability (Helstrom bound) is given by [10, 6]:
$P_{\text{HB}}^{\text{DSS}}\left(\alpha,r\right)$ $=\frac{1}{2}\left(1-\sqrt{1-\left|\left\langle\psi_{0}|\psi_{1}\right\rangle\right|^{2}}\right)$ (2)
@@
which $W\left(x,p|\rho_{i}\right)$ is the Wigner function of $\rho_{i}$ (see Appendix). And the resulting error probability is
$P_{\text{SQL}}^{\text{DSS}}=\int_{0}^{+\infty}P\left(x|\rho_{0}\right)\mathrm{d}x=\frac{1}{2}\text{erfc}\left(\sqrt{2}\alpha\text{e}^{r}\right)$ (4)
@@
$\beta:=\frac{\sinh^{2}(r)}{N}$ (5)
@@
Accordingly, both the Helstrom bound and the SQL for S-BPSK become functions of $\left(N,\beta\right)$, denoted $P_{\text{SQL}}^{\text{DSS}}\left(N,\beta\right)$, and $P_{\text{HB}}^{\text{DSS}}\left(N,\beta\right)$. The squeezing fraction can be optimized at fixed total energy. By minimizing $P_{\text{HB}}^{\text{DSS}}\left(N,\beta\right)$ over $\beta$ one obtains the optimal value [6]:
$\beta_{\text{opt}}(N)=\frac{N}{2N+1}$ (7)
Therefore, $\alpha^{2}\text{e}^{2r}=N(N+1)$. Both the Helstrom bound and the SQL for S-BPSK are rewritten as:
@@
$P_{\text{SQL}}^{\text{DSS}}\left(N,\beta_{\text{opt}}\right)=\frac{1}{2}\text{erfc}\left(\sqrt{2N\left(N+1\right)}\right)$ (8)
$P_{\text{HB}}^{\text{DSS}}\left(N,\beta_{\text{opt}}\right)=\frac{1}{2}\left(1-\sqrt{1-\exp\left\{-4N(N+1)\right\}}\right)$ (9)
@@
Figure 1: The ratios of $P_{\text{HB}}^{\text{DSS}}\left(N,\beta_{\text{opt}}\right)$, $P_{\text{SQL}}^{\text{DSS}}\left(N,\beta_{\text{opt}}\right)$ and $P_{\text{SQL}}^{\text{CS}}\left(N\right)$ to $P_{\text{HB}}^{\text{CS}}\left(N\right)$ vs. input state energy N. Here, $P_{\text{SQL}}^{\text{CS}}\left(N\right)=\frac{1}{2}\left[1-\mathrm{erf}\left(\sqrt{2\cdot N}\right)\right]$ and $P_{\text{HB}}^{\text{CS}}\left(N\right)=\frac{1}{2}\left[1-\sqrt{1-\exp\left(-4\cdot N\right)}\right]$.
@@
When the DSSs are prepared with this choice, the S-BPSK Helstrom bound is significantly lower than that of C-BPSK at the same energy. In particular, for $N\approx 1$ the S-BPSK Helstrom bound already reaches the order of $10^{-4}$, whereas C-BPSK requires approximately twice the energy to achieve a comparable error probability. Moreover, in the low-photon number regime the S-BPSK SQL can fall below the Helstrom bound of C-BPSK, with the crossover occurring around $N\approx 0.659$ (see Fig. 1). At $N=1.0$, the HB for S-BPSK surpasses the SQL for C-BPSK by a remarkable 17.39dB, and the SQL for S-BPSK also is 2.94dB lower. In the subsequent analyses, all DSSs employ the optimal squeezing fraction $\beta_{\text{opt}}$.
III Quantum-enhanced Receiver for discriminating BPSK Displaced Squeezed Vacuum States
We propose a quantum-enhanced receiver for S-BPSK discrimination, referred to as the displacement–squeeze receiver (DSR). Its schematic is shown in Fig. 2. Compared with conventional displacement receivers, the DSR introduces an additional squeezing operation. The squeezing strength matches that used at the transmitter, whereas the squeezing axis is rotated by $\pi/2$ in phase space.
Figure 2: Schematic of the proposed DSR. The input state $\rho_{i}$ is transformed to $\zeta_{i}$ via a displacement followed by a squeezing operation, $U=S(-r)D(\alpha)$, and is then measured by a photon-number-resolving (PNR) detector with outcome $n$. The final decision is made according to the maximum a posteriori (MAP) criterion.
In the DSR, the input state $\rho_{i}=\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|$ first interferes with a local oscillator $\left|\alpha\right\rangle$ on a highly transmissive beam splitter ($\tau\to 1$), implementing the displacement operation $D(\alpha)$. This maps the BPSK pair $\left\{\left|-\alpha,r\right\rangle,\left|+\alpha,r\right\rangle\right\}$ to an OOK-like pair $\left\{\left|0,r\right\rangle,\left|+2\alpha,r\right\rangle\right\}$. Subsequently, a squeezing operation S(r)S(-r) is applied, squeezing the $P$ quadrature. The processed state is denoted by
$\zeta_{i}=U\rho_{i}U^{\dagger},\quad U=S(-r)D(\alpha).$ (10)
A PNR detector then measures $\zeta_{i}$ and outputs the photon number $n$, based on which the receiver makes a MAP decision.
@@ III Quantum-enhanced Receiver for discriminating BPSK Displaced Squeezed Vacuum States
Figure 3: Phase-space representations (a–c) and Fock-basis populations (d–f) of the input states $\rho_{i}$ in the DSR for mean photon number $N=1.0$ and squeezing fraction $\beta=\beta_{\text{opt}}$. Panels (a) and (d) show the input states $\rho_{i}$; panels (b) and (e) show the states after the displacement operation $D\rho_{i}D^{\dagger}$; and panels (c) and (f) show the states after the full DSR transformation $\zeta_{i}=U\rho_{i}U^{\dagger}$.
We next provide an intuitive explanation of the DSR advantage. Figs. 3(a)–(c) illustrate how the phase-space distributions evolve under the displacement and squeezing operations for $N=1.0$ and $\beta=\beta_{\text{opt}}$. The two input states $\rho_{0}$ and $\rho_{1}$ are separated by $\Delta=2\alpha$ in phase space. After the displacement operation, the two states become a squeezed vacuum state and a displaced squeezed vacuum state, respectively (Fig. 3(b)):
$\left\{\begin{aligned} &D(\alpha)\cdot\left|\psi_{0}\right\rangle=S(r)\left|0\right\rangle\\ &D(\alpha)\cdot\left|\psi_{1}\right\rangle=D(2\alpha)S(r)\left|0\right\rangle\end{aligned}\right$. (11)
The displacement does not change the separation $\Delta=2\alpha$ between the state centroids in phase space; however, it leads to distinct Fock-basis populations for $\rho_{0}$ and $\rho_{1}$ (see Figs. 3(d) and (e)).
After the subsequent squeezing operation, the two states evolve into the vacuum state and a coherent state, respectively (Fig. 3(c)):
$\left\{\begin{aligned} &U\cdot\left|\psi_{0}\right\rangle=\left|0\right\rangle\\ &U\cdot\left|\psi_{1}\right\rangle=S(-r)D(2\alpha)S(r)\left|0\right\rangle=D(\gamma)\left|0\right\rangle=\left|\gamma\right\rangle\end{aligned}\right$. (12)
Here,
$\gamma$ $=2\alpha\cdot\cosh(r)+2\alpha^{*}\cdot\sinh(r)$ (13)
$=2\alpha\cdot\mathrm{e}^{r}>2\alpha,$
where the last equality assumes $\alpha$ is real (i.e., the phase reference is chosen such that $\alpha\in\mathbb{R}$). Therefore, the S-BPSK input $\rho_{i}\in\left\{\left|-\alpha,r\right\rangle,\left|+\alpha,r\right\rangle\right\}$ is converted into a coherent-state OOK pair $\zeta_{i}\in\left\{\left|0\right\rangle,\left|\gamma\right\rangle\right\}$. Importantly, the phase-space separation increases from $2\alpha$ to $\gamma=2\alpha\,\mathrm{e}^{r}$, enhancing distinguishability for subsequent PNR detection. By comparison, a coherent-state displacement receiver (without squeezing) yields a separation $\Delta_{\mathrm{C}}=2\sqrt{N}$, whereas the DSR produces $\gamma=2\sqrt{N(N+1)}$.
Figs. 3(d)–(f) show the corresponding changes in Fock-basis populations. After the squeezing operation (Fig. 3(f)), the population associated with $\rho_{1}$ shifts toward higher photon numbers, while the population associated with $\rho_{0}$ concentrates near $n=0$. Consequently, the overlap between the two photon-number distributions is reduced, facilitating discrimination with a PNR detector.
III.1 PNR detection model and MAP decision rule
We now incorporate finite PNR resolution. A PNR detector with resolution $M$ is denoted as PNR(M)(M) and is described by the POVM elements
$\Pi_{n}=\left|n\right\rangle\left\langle n\right|,\quad n=0,1,.,M-1$ (14)
$\Pi_{M}=\mathbf{I}-\sum_{i=0}^{M-1}\Pi_{i}$
where $\mathbf{I}$ is the identity operator in the Fock basis. For a measured coherent state $\rho$ with mean photon number $\mu$, the conditional probability of obtaining outcome $n$ is
$q_{n}^{M}\left(\mu\right)=\mathbf{Tr}\left\{\Pi_{n}\cdot\rho\right\}=\left\{\begin{aligned} &\mathrm{e}^{-\mu}\cdot\frac{\mu^{n}}{n!},\quad n\leq M-1\\ &1-\mathrm{e}^{-\mu}\cdot\sum_{j=0}^{M-1}\frac{\mu^{j}}{j!},\quad n=M\end{aligned}\right$. (15)
Accordingly, the conditional probability of outcome $n$ given that symbol $i$ was sent is
@@ III.1 PNR detection model and MAP decision rule
$P(n|i)=\mathbf{Tr}\left(\Pi_{n}\cdot\zeta_{i}\right)=\left\{\begin{aligned} &q_{n}^{M}(0),\ i=0\\ &q_{n}^{M}\left(|\gamma|^{2}\right),\ i=1\end{aligned}\right$. (16)
The MAP decision rule for binary discrimination can be equivalently expressed as an optimal threshold test on $n$ [28, 20], summarized in Table 1. The corresponding threshold $n_{\text{th}}^{*}$ satisfies Eq. (17).
$P_{0}\cdot P(n|0)\gtreqless P_{1}\cdot P(n|1),\ \text{where}\ n\lesseqgtr n_{\text{th}}^{*}$ (17)
$\Rightarrow n_{th}^{*}=\min\left\{\lceil n\rceil,\ \text{M}\right\}$
where, $\lceil*\rceil$ denotes the ceiling function (rounding up); Under ideal and equal-prior probability conditions, one obtains $n_{\text{th}}^{*}=1$.
Table 1: Threshold decision strategy of the DSR. If the PNR(M)(M) outcome satisfies $n<n_{\text{th}}^{*}$, decide symbol ‘0’; otherwise decide symbol ‘1’.
The outcomes of PNR(M) Decision
$n\geq n_{\text{th}}^{*}$ symbol ’1’
$n<n_{\text{th}}^{*}$ symbol ’0’
@@
III.2 Error probability under ideal conditions
Now we analyze the error probability of the DSR. Assuming the optimal threshold value is $n_{th}^{*}$, the threshold decision strategy (see Table. 1) can be expressed using POVM operators:
$\text{symbol}\ ^{\prime}0^{\prime}:\quad\Pi_{\text{off}}=\sum_{n=0}^{n_{th}^{*}-1}\Pi_{n}=\sum_{n=0}^{n_{th}^{*}-1}\left|n\right\rangle\left\langle n\right|$ (18)
$\text{symbol}\ ^{\prime}1^{\prime}:\quad\Pi_{\text{on}}=\mathbf{I}-\Pi_{\text{off}}$
The average error probability is then
$P_{\text{err}}^{\text{DSR}}$ $=1-\frac{1}{2}\left[P(\text{C}|0)+P(\text{C}|1)\right]$ (19)
$=1-\frac{1}{2}\left[\mathbf{Tr}\left(\Pi_{\text{off}}\cdot\zeta_{0}\right)+\mathbf{Tr}\left(\Pi_{\text{on}}\cdot\zeta_{1}\right)\right]$
where $P(\text{C}|i)$ denotes the probability of a correct decision conditioned on sending symbol $i$. Under ideal conditions, $n_{\text{th}}^{*}=1$, hence $\Pi_{\text{off}}=\left|0\right\rangle\left\langle 0\right|,\Pi_{\text{on}}=\mathbf{I}-\left|0\right\rangle\left\langle 0\right|$. The DSR error probability becomes
$P_{\text{err}}^{\text{DSR}}$ $=1-\frac{1}{2}\left[q_{0}^{M}(0)+1-q_{0}^{M}\left(|\gamma|^{2}\right)\right]$ (20)
$=\frac{1}{2}\exp\left\{-4|\alpha|^{2}\mathrm{e}^{2r}\right\}=\frac{1}{2}\exp\left\{-4N\left(N+1\right)\right\}$
Eq. (20) shows that, under ideal conditions, $P_{\text{err}}^{\text{DSR}}$ is independent of the PNR resolution $M$. Therefore, in the numerical results of Fig. 4, we consider PNR(1), i.e., a single-photon detector (SPD).

### v2

# Near-optimal discrimination of displaced squeezed binary signals using displacement, inverse-squeezing, and photon-number-resolving detection

@@ Abstract
We propose an inverse-squeezing Kennedy receiver for discriminating binary phase-shift-keyed displaced squeezed vacuum states. The receiver combines a Kennedy-type nulling displacement, an orthogonally oriented inverse-squeezing operation and photon-number-resolving detection with a maximum-a-posteriori threshold rule. Its key mechanism is that the inverse-squeezing stage converts transmitter-side squeezing into enhanced photon-number contrast, or equivalently an effective coherent-state separation gain, that can be directly exploited at the measurement stage. Under ideal equal-prior conditions, the receiver surpasses the standard quantum limit for squeezed-state binary phase-shift keying at approximately $N\approx 0.3$, outperforms the Helstrom bound of coherent-state binary phase-shift keying at approximately $N\approx 0.4$, and reaches the 1% error level near $N\approx 0.6$. We further analyze its performance under realistic imperfections, including finite detector efficiency, dark counts, channel phase diffusion, receiver thermal noise and transmission loss. The results show that adaptive thresholding preserves robust performance against detector and noise imperfections over practical parameter ranges, whereas transmission loss progressively suppresses the squeezing-enabled advantage. These findings indicate that, for the fixed source parametrization adopted in this work, the proposed receiver is most advantageous in the low-loss regime, especially at low source energies.
@@ I Introduction
Quantum state discrimination is a central problem in quantum information and optical communications, because the minimum achievable detection error directly determines the reliability of information transfer. Among optical encodings, coherent states have long played a dominant role owing to their experimental accessibility and their robustness under bosonic loss [15, 38, 23, 37]. This has led to the development of a wide range of receiver architectures for coherent-state discrimination, including the Kennedy receiver[22], optimized-displacement receivers[34], multistage and feedforward receivers[35, 21, 7, 25, 42], hybrid detection strategies [24, 1, 27, 43]. Nevertheless, coherent states remain non-orthogonal, and hence their discrimination error probability is fundamentally bounded by the Helstrom bound (HB) [17, 18].
In parallel with these classical developments, recent years have witnessed substantial progress in coherent-state discrimination beyond the conventional binary signal. These advances include photon-counting and structured optical receivers approaching the quantum limit [12, 31], receiver designs for multi-symbol and multi-mode constellations [2, 33, 20, 30, 32], and learning-assisted strategies for receiver optimization, calibration, and near-optimal joint detection [29, 11, 4]. Such studies have significantly broadened the scope of quantum receiver theory, both in terms of signal alphabets and measurement strategies. However, most of these developments still focus on coherent-state alphabets, i.e., on improving the receiver side while keeping the transmitted states classical in the phase-space sense.
A different route is to exploit nonclassical transmitter states whose intrinsic overlap is smaller than that of coherent states at the same energy [28, 10, 36]. In this context, Ref. [10] showed that squeezed-vacuum-based phase-shift-keyed communication can outperform coherent-state binary phase-shift keying (BPSK) benchmarks. In particular, by displacing a squeezed vacuum state with opposite phases, one obtains the binary alphabet of phase-opposite displaced squeezed vacuum states,
@@
which we refer to as squeezed-state BPSK (S-BPSK). For a fixed mean signal energy, the Helstrom bound of S-BPSK can be significantly lower than that of coherent-state BPSK (C-BPSK), and even the homodyne limit of S-BPSK may fall below the Helstrom bound of C-BPSK in the low-energy regime [10]. These results indicate that transmitter-side squeezing can serve as a genuine communication resource for error reduction.
Motivated by this observation, the key open question is how to design a practical receiver that can efficiently convert the squeezing resource of the transmitted S-BPSK signals into an experimentally accessible discrimination advantage. In this work, we address this problem by proposing the inverse-squeezing Kennedy receiver (IS-Kennedy), which combines a displacement operation, an orthogonally oriented inverse-squeezing (IS) operation, and photon-number-resolving (PNR) detection with a maximum-a-posteriori (MAP) threshold rule. The basic idea is to map the incoming displaced squeezed-state alphabet into a more distinguishable effective coherent-state on–off-keying pair, thereby increasing the phase-space separation and reducing the overlap of the corresponding photon-number statistics. Under ideal conditions, we show that the error probability of the proposed receiver always remains within a factor of two of the S-BPSK Helstrom bound for equal priors, while outperforming coherent-state benchmarks in the low-energy regime. We further analyze its robustness under realistic imperfections, including finite detector efficiency, dark counts, phase diffusion, and receiver thermal noise.
The structure of this paper is as follows: In Sec. II, we review the signal model and benchmark limits for S-BPSK discrimination. In Sec. III, we introduce the IS-Kennedy receiver and analyze its operation and performance under ideal conditions. In Sec. IV, we study its performance in the presence of nonideal detectors and environmental noise. Finally, Sec. V concludes the paper.
@@ II Channel model and performance benchmarks
In this section we summarize the signal model and the fundamental performance limits for BPSK communication based on displaced squeezed vacuum states, and we fix the notation used in the rest of the paper.
For binary discrimination between quantum states $\left|\psi_{0}\right\rangle$ and $\left|\psi_{1}\right\rangle$ with prior probabilities $\pi_{0}=\pi_{1}=\frac{1}{2}$, the minimum error probability (Helstrom bound) is given by [17, 10, 5]:
$P_{\text{HB}}^{\text{S-BPSK}}\left(\alpha,r\right)$ $=\frac{1}{2}\left(1-\sqrt{1-\left|\left\langle\psi_{0}|\psi_{1}\right\rangle\right|^{2}}\right)$ (2)
@@
which $W\left(x,p|\rho_{i}\right)$ is the Wigner function of $\rho_{i}$ (see Appendix A). And the resulting error probability is
$P_{\text{SQL}}^{\text{S-BPSK}}=\int_{0}^{+\infty}P\left(x|\rho_{0}\right)\mathrm{d}x=\frac{1}{2}\text{erfc}\left(\sqrt{2}\alpha\text{e}^{r}\right)$ (4)
@@
$\beta\triangleq\frac{\sinh^{2}(r)}{N}$ (5)
@@
Accordingly, both the Helstrom bound and the SQL for S-BPSK become functions of $\left(N,\beta\right)$, denoted $P_{\text{SQL}}^{\text{S-BPSK}}\left(N,\beta\right)$, and $P_{\text{HB}}^{\text{S-BPSK}}\left(N,\beta\right)$. The squeezing fraction can be optimized at fixed total energy. By minimizing $P_{\text{HB}}^{\text{S-BPSK}}\left(N,\beta\right)$ over $\beta$ one obtains the optimal value [10]:
$\beta_{\text{opt}}(N)=\frac{N}{2N+1},\quad\alpha\text{e}^{r}=\sqrt{N(N+1)}$ (7)
Both the Helstrom bound and the SQL for S-BPSK are rewritten as:
@@
$P_{\text{SQL}}^{\text{S-BPSK}}\left(N,\beta_{\text{opt}}\right)=\frac{1}{2}\text{erfc}\left(\sqrt{2N\left(N+1\right)}\right)$ (8)
$P_{\text{HB}}^{\text{S-BPSK}}\left(N,\beta_{\text{opt}}\right)=\frac{1}{2}\left(1-\sqrt{1-\exp\left\{-4N(N+1)\right\}}\right)$ (9)
@@
Figure 1: The ratios of $P_{\text{HB}}^{\text{S-BPSK}}\left(N,\beta_{\text{opt}}\right)$, $P_{\text{SQL}}^{\text{S-BPSK}}\left(N,\beta_{\text{opt}}\right)$ and $P_{\text{SQL}}^{\text{C-BPSK}}\left(N\right)$ to $P_{\text{HB}}^{\text{C-BPSK}}\left(N\right)$ vs. input state energy N. Here, $P_{\text{SQL}}^{\text{C-BPSK}}\left(N\right)=\frac{1}{2}\left[1-\mathrm{erf}\left(\sqrt{2\cdot N}\right)\right]$ and $P_{\text{HB}}^{\text{C-BPSK}}\left(N\right)=\frac{1}{2}\left[1-\sqrt{1-\exp\left(-4\cdot N\right)}\right]$.
@@
When the S-BPSK are prepared with this choice, the S-BPSK Helstrom bound is significantly lower than that of C-BPSK at the same energy. In particular, for $N\approx 1$ the S-BPSK Helstrom bound already reaches the order of $10^{-4}$, whereas C-BPSK requires approximately twice the energy to achieve a comparable error probability. Moreover, in the low-photon number regime the S-BPSK SQL can fall below the Helstrom bound of C-BPSK, with the crossover occurring around $N\approx 0.659$ (see Fig. 1). At $N=1.0$, the Helstrom bound for S-BPSK surpasses the SQL for C-BPSK by a remarkable 17.39dB, and the SQL for S-BPSK also is 2.94dB lower. In the subsequent analyses, all S-BPSK employ the optimal squeezing fraction $\beta_{\text{opt}}$.
III inverse-squeezing Kennedy receiver under ideal conditions
In this section, we propose a novel quantum-enhanced receiver for S-BPSK discrimination, referred to as the IS-Kennedy. Its schematic is shown in Fig. 2. Compared to the conventional Kennedy receiver, the IS-Kennedy incorporates two key modifications: (i) the insertion of an IS module after the displacement operation, which applies the operator S(r)S(-r) to the signal; and (ii) the upgrade of the terminal single-photon detector (SPD) to a photon-number-resolving (PNR) detector, enabling MAP decision-making under generalized statistical models. The squeezing amplitude of the IS module is matched to that of the transmitter, while the squeezing axis is rotated by $\pi/2$ (orthogonal) relative to the transmitter. This configuration cancels the initial squeezing and produces an equivalent amplification of the signal’s mean photon number. Effectively, the IS-Kennedy converts the transmitter’s squeezing resources into an equivalent amplitude gain at the receiver, thereby reducing the S-BPSK discrimination problem to an “amplified Kennedy nulling problem”. Furthermore, the PNR detector provides the necessary statistical degrees of freedom for optimal MAP decisions under non-ideal conditions, which will be discussed later.
[…]
Subsequently, $\sigma_{i}$ passes through the IS module, mapping the S-OOK alphabet to a coherent-state on-off keying (C-OOK) alphabet $\zeta_{i}$:
[…]
In the general case, $\gamma=\alpha\cosh r+\alpha^{*}\sinh r$. Under the condition $\alpha\in\mathbb{R}_{+}$ adopted in this work, $\gamma=\alpha e^{r}$. Therefore, we define:
$\gamma\triangleq\alpha e^{r},\qquad N_{\text{eff}}\triangleq|\gamma|^{2}=|\alpha|^{2}e^{2r}.$ (12)
With the optimal energy allocation $\beta_{\text{opt}}$, the effective photon number can be further expressed as $N_{\text{eff}}=N(N+1)>N$. This indicates that the IS module transforms the transmitter’s squeezing resources into a “coherent-state separation gain” directly accessible by the receiver. This results in a larger Euclidean distance between the two states in phase space (Fig. 3(a-c)) and simultaneously reduces the overlap of the Fock-basis populations of $\zeta_{i}$ (Fig. 3(e-f)).
@@
Figure 3: Phase-space representations (a–c) and Fock-basis populations (d–f) of the input states $\rho_{i}$ in the IS-Kennedy for mean photon number $N=1.0$ and squeezing fraction $\beta=\beta_{\text{opt}}$. Panels (a) and (d) show the input states $\rho_{i}$; panels (b) and (e) show the states after the nulling displacement $\sigma_{i}=D(\alpha)\rho_{i}D^{\dagger}(\alpha)$. Panels (c) and (f) show the states after inverse-squeezing $\zeta_{i}=U\rho_{i}U^{\dagger}$ with $U=S(-r)D(\alpha)$.
Next, PNR detection is performed on $\zeta_{i}$, yielding an output photon number $n$. The positive operator-valued measure (POVM) for an ideal PNR detector is given by:
[…]
From Eq. (11), we have $|\zeta_{0}=\ket{0}\bra{0}$ and $|\zeta_{1}=\ket{2\gamma}\bra{2\gamma}$, with mean photon numbers $\mu_{0}=0$ and $\mu_{1}=|2\gamma|^{2}=4N_{\text{eff}}$, respectively. Thus, the probability of detecting $n$ photons given transmitted symbol $i$ is:
@@
$P(n|i)=\text{Tr}\left(\zeta_{i}\Pi_{n}^{\text{ideal}}\right)=\left\{\begin{aligned} &\text{Poiss}\left(n;\mu_{0}\right)=\delta_{0,n},&i=0,\\ &\text{Poiss}\left(n;\mu_{1}\right),&i=1,\end{aligned}\right$. (14)
where $\text{Poiss}\left(n;\mu\right)=\mathrm{e}^{-\mu}\frac{\mu^{n}}{n!}$.
The posterior probabilities are given by:
$P(i=0|n)=\frac{\pi_{0}P(n|i=0)}{\sum_{k}\pi_{k}P(n|i=k)},\quad P(i=1|n)=\frac{\pi_{1}P(n|i=1)}{\sum_{k}\pi_{k}P(n|i=k)}.$ (15)
The MAP decision rule is defined as:
$\hat{i}(n)=\begin{cases}1,&P(i=1|n)\geq P(i=0|n),\\ 0,&\text{otherwise}.\end{cases}$ (16)
where $\hat{i}$ is the decision of IS-Kennedy receiver. Assuming equal priors ($\pi_{0}=\pi_{1}=1/2$), the log-likelihood ratio is:
$L(n)=\ln\frac{P(n|i=0)}{P(n|i=1)}=\ln\frac{\text{Poiss}\left(n;\mu_{0}\right)}{\text{Poiss}\left(n;\mu_{1}\right)}=-\mu_{0}+\mu_{1}+n\ln\left(\frac{\mu_{0}}{\mu_{1}}\right).$ (17)
Since $L(n)$ is a monotonic function of $n$, the MAP decision (Eq. (16)) is equivalent to a threshold detection:
$\hat{i}(n)=\begin{cases}1,&n\geq n_{\text{th}}^{*},\\ 0,&\text{otherwise}.\end{cases}$ (18)
The optimal threshold $n_{\text{th}}^{*}$ corresponds to the zero-crossing of the log-likelihood ratio:
$n_{\text{th}}^{*}=\left\lceil\frac{\mu_{1}-\mu_{0}}{\ln\mu_{1}-\ln\mu_{0}}\right\rceil.$ (19)
Under ideal conditions where $\mu_{0}=0$, we have $\lim_{\mu_{0}\to 0}\frac{\mu_{1}-\mu_{0}}{\ln\mu_{1}-\ln\mu_{0}}\to 0^{+}$; thus, $n_{\text{th}}^{*}=1$. This corresponds to a standard on-off detection strategy, implying that a simple SPD suffices to achieve PNR performance in this ideal condition. It is important to emphasize, however, that the introduction of the PNR detector is not intended to improve the performance of the IS-Kennedy under ideal condition. Rather, it addresses practical system imperfections: once detector imperfections (efficiency/dark counts) or environmental noise occur, $\zeta_{0}$ deviates from the vacuum state, yielding a non-zero $P(n\geq 1|i=0)$. In such cases, the optimal decision is no longer a simple on-off rule, and the PNR detector provides the necessary statistical information to implement a generalized MAP or optimal threshold strategy. Quantitative analysis of these non-ideal factors is presented in subsequent sections.
The false alarm probability $P_{\text{FA}}$ (deciding 1 when 0 is sent) and the miss detection probability $P_{\text{Mi}}$ (deciding 0 when 1 is sent) are:
$P_{\text{FA}}=P\left(n\geq n_{\text{th}}^{*}\mid i=0\right),\quad P_{\text{Mi}}=P\left(n<n_{\text{th}}^{*}\mid i=1\right).$ (20)
@@
Therefore, under equal priors and ideal conditions ($n_{\text{th}}^{*}=1$), the average error probability of the IS-Kennedy is:
$P_{\text{err}}^{\text{ISK},\text{ideal}}$ $=\frac{1}{2}\left(P_{\text{FA}}+P_{\text{Mi}}\right)$ (21)
$=\frac{1}{2}\left[1-P(n=0|i=0)+P(n=0|i=1)\right]$
$=\frac{1}{2}\exp\left(-|2\gamma|^{2}\right)=\frac{1}{2}\exp\left(-4N_{\text{eff}}\right).$
Substituting the optimal energy allocation condition $N_{\text{eff}}=N(N+1)$, we obtain:
$P_{\text{err}}^{\text{ISK},\text{ideal}}=\frac{1}{2}\exp\left[-4N(N+1)\right].$ (22)
Notably, we observe that:
$P_{\text{err}}^{\text{ISK},\text{ideal}}(N)=P_{\text{err}}^{\text{K},\text{ideal}}(N_{\text{eff}}),$ (23)
where $P_{\text{err}}^{\text{K},\text{ideal}}$ is the error probability of the conventional Kennedy receiver under ideal conditions [22]. This implies that, under matched squeezing, the IS-Kennedy is equivalent to mapping an S-BPSK signal with mean photon number $N$ to a C-BPSK signal with mean photon number $N_{\text{eff}}$ via the IS module, followed by Kennedy nulling and on-off detection.